# A Survey of Passive Sensing in the Workplace

### v2

@@ IV-B PRODUCTIVITY
Therefore, researchers in the pervasive computing community have often turned to job performance inventories developed by psychologists as ground truth to measure perceived workplace performance across organizations and industries in a generalizable manner. They have used passive sensing to predict participant scores for these inventories, or to classify individuals into higher and lower performers within the workplace. The inventories usually measure either task performance such as Individual Task Proficiency (ITP), In-Role Behavior (IRB); or assess behaviors that promote the effectiveness of organizations and its members such as Interpersonal and Organizational Deviance (IOD-ID, IOD-OD), Organizational Citizenship Behavior (OCB).
@@
Task performance refers to duties or actions that are formally recognized and rewarded by management [51]. Several studies attempt to predict ITP and IRB as measures of task performance. In a study of 84 full-time nursing professionals over a ten-week period, Feng et al. [32] use Fitbit and Bluetooth proximity data to study how an individual’s physiological responses vary according to the work environment (room and function of the room). The authors compute a measure of mutual dependency of time in each room in the hospital and physiological responses and find this measure along with baseline aggregated features (e.g., min, mean, max, SD, etc.) leads to predictions of ITP with an adjusted R2 score of .11 and .12 for IRB. Another study [34] recruits 67 nurses and instruments them with a sociometric badge capable of measuring physical activity, speech activity, face to face interaction and social network inferred through the sociometric badge proximity sensor. Multiple linear regression models with sensed features as predictors are able to predict 49% of the variance in groups’ perception of workload as measured in an ad-hoc survey and 63% of the variance in groups’ perception of productivity.
In studies of information workers, several works stand out. The study by Das Swain et al. [33] recruits 249 individuals over 62 days and used Bluetooth beacons to analyze desk and away-from-desk sessions per hour, time at work and time at home. The authors produce a measure of organization fit by analyzing the convergence of individual routines in the aforementioned behaviors and their organization’s normative behavior. They find that a linear model with Big Five Personality [52] measures along with routine fit could predict up to 28% of the variance in IRB, with routine fit having a significant (p <<< 0.05) positive association with IRB. In another study of 603 information workers over 110 days, Das Swain et al. [35] propose the concept of organizational personas, a clustering-based approach in which personality facets are constructed over Big Five Personality data and activity facets are constructed from multi-modal data from a fitness tracker, a phone agent (location, phone usage) and Bluetooth beacons (desk sessions). The authors report that the activity facets account for a small but statistically significant part of the variance in ITP. In a similar study of 554 information workers using a phone agent, beacons and a Garmin fitness tracker, Mirjafari et al. [4] use K-Means clustering on the participant averages of ITP, IRB, OCB, IOD-ID and IOD-OD scores to classify them into higher and lower performers in their organizations. Results indicate that features from the fitness tracker alone can discriminate performers with an AUROC of .72; features from the phone agent can discriminate performers with an AUROC of .65 and features from both can do so with an AUROC of .83; showing how multi-modality can better explain these complex behaviors.
In a follow-up study of 298 information workers, Mirjafari et al. [53] used auto-encoder-generated features based on passive sensing data from mobile phones and a Garmin fitness tracker to predict day-to-day job performance dynamics, i.e., detecting whether there was an improvement, decline, or no change in the job performance for individuals in the study as assessed by ITP, IRB, OCB, IOD-ID and IOD-OD. The final model used XGBoost to achieve an F1 score of 75%, surpassing the authors’ baseline of 33%. The authors also show the use of Gradient Analysis, a technique popular in the computer vision domain [54, 55], to provide user-actionable interpretation of the model, demonstrating the potential to inform individuals of what behaviors should be modified in order to improve their performance. Similarly, Mark et al. [29] captured time spent on email interactions of 40 information workers using an activity logging software finding that the more time people spent on emails daily, the lower their reported levels of productivity is.
In the behavioral side of workplace performance inventories, IOD-ID and IOD-OD are measures of “bad” conduct in the workplace. Behaviors that indicate IOD-ID can involve cursing a co-worker, playing pranks, or making fun of someone. Behaviors that indicate IOD-OD can be tardiness or absenteeism, leaving work early without permission, putting little effort into work, among others [56]. On the other hand, OCB is a measure of “good” behavior. OCB is composed of behaviors that reflect altruism, conscientiousness, sportsmanship, courtesy and civic virtue [57]. Several studies have focused on these measures.
Feng and Narayanan [10] propose a method for capturing behavioral consistency in wearable data using the activity curve model. They find that consistency features improve accuracy by up to 6% when compared to using only summary features from the Fitbit fitness tracker in a study of 97 hospital workers throughout 10 weeks. Nepal et al. [36] use commute based passive sensing data (viz. location-related context, physiology, variability) from phones and wearables to predict the OCB and IOD score of workers. Authors collect passive data of 275 workers throughout a year while the workers are commuting to and from work. They train a stacked machine learning model on the commute data and achieve MAEs of less than 10% of each performance ground truths, suggesting that passive sensors may be capable of measuring an individual’s commuting experience and its impact on their job performance. In a previously mentioned study by Feng et al. [32], the mutual dependency between physiological response and time spent in different environments at the workplace improves the prediction of IOD-ID (adjusted R2 = 0.065) and IOD-OD (adjusted R2 = 0.092) as well. Mirjafari et al. [4] find in their classification of higher and lower performers that passive sensing data is associated with IOD-ID, IOD-OD and OCB. Das Swain et al. [35] also report that the activity facet created from passively sensed data is associated with IOD-OD, IOD-ID and OCB. They also find that a better fit between individual and organizational desk routines measured through beacon proximity implied lower IOD [33]. However, maximum accuracy was still low with 62.4% for IOD-ID, 60.6% for IOD-OD, 56.6% for OCB, and 57.8% for STAI.
The work by Nepal et al. [37] follows a different line of research. Instead of attempting to predict subjective performance measures from objective data, it intends to find what objective data is related to individual success in organizations, i.e., getting promoted. The authors note that if promotion can be detected and it produces objective physiological signals, researchers could learn from these signals and through the study of job promotion – which could be inferred – what behaviors could be behind the high performance that granted the promotion to measure them objectively, foster them and replicate them.
@@
In the modern era, nurturing collaboration through open offices has become the norm and workers face necessary interruptions throughout their workday to efficiently tackle problems. However, there are studies demonstrating that the nature and timing of interruptions negatively affect workplace productivity [58, 59] and increase worker stress [60]. Therefore, evaluating interruptibility has significant value to companies in terms of cost and time. To continuously measure a person’s interruptibility, Zuger et al. [39] study 13 software engineers over two weeks using wearable sensors in addition to keyboard and mouse interaction data. The ground truth in this study was collected using a pop-up question during work which asked the user to rate their interruptibility on a 1-7 Likert scale, and these ratings are further grouped into three states (12—345—67) for classification. Ultimately, they predicted interruptibility accurately (68.3%) by training a random forest on several features extracted from sensor modalities like heart rate, activity, and sleep.
Other works target efforts to support the design of future interventions in the workplace. Kimani et al. [30] created a conversational agent designed to assist information workers in achieving various work-related objectives, such as task scheduling and prioritization, task switching, providing reminders to take breaks, minimizing social media distractions, and reflecting on completed tasks. This agent also harnessed contextual data about the user to determine the optimal times for proactive intervention. Implemented among 24 information workers for a duration of six days, the agent analyzed speech, facial imagery, and computer usage patterns. The findings indicated that many participants not only appreciated the agent’s recommendations regarding breaks and work reflections but also became more conscious of their work habits. This heightened awareness led some to alter their routines for enhanced productivity and wellbeing. Nepal et al.[31] develop a Large Language Model (LLM) powered personalized productivity agent that utilizes computer-based telemetry data from information workers to provide tailored assistance. Authors highlight the importance of user-centric design, adaptability, and the balance between personalization and privacy in AI-assisted productivity tools. Similarly, Kadoya et al. [61] use a wearable to derive control measures to understand the relation of emotional states and productivity in the workplace. Such understanding would help in the future to focus interventions that can improve the emotional state of workers. Mark et al. [22] found that affect balance correlates with concentration and workplace productivity, and sleep measured by a fitness tracker was related to affect balance. In another study, Schaule et al. [38] designed a proof of concept system that integrated a smartwatch to predict office workers’ cognitive load and suggest when users were available to be interrupted. Similarly, Di Lascio et al. [40] show that wearables can predict workplace activity passively, which has the potential to be used as the foundation of systems to automatically block distractions during focus hours and to suggest breaks. Kucukozer-Cavdar et al. [41] also create a model for predicting office workers’ availability and inclination to take breaks using mobile sensing features relevant to workers’ location, activity, application usage, ringer mode. Das Swain et al.[43] evaluated the effect of computer-assisted protected (CAP) time on 89 information workers. Through a randomized control trial in naturalistic conditions, they found that CAP significantly benefits workers who have never used such automated schedulers. Similarly, Saha et al.[44] found that CAP improved positive well-being factors such as relaxation while reducing negative aspects like anger. To summarize, these studies not only advance our understanding of workplace performance and the behaviors that bolster productivity but also showcase the pervasive computing community’s active efforts in crafting interventions that enhance performance. With the rapid advancement and integration of LLMs and artificial intelligence (AI), this field stands on the brink of a significant growth trajectory. The combination of passive sensing with the nuanced analytical capabilities of LLMs presents an exciting next frontier, promising to redefine and amplify intervention strategies for workplace productivity and wellbeing.
V DISCUSSION
@@
V-A Considerations for Ground Truth
@@
V-B Considerations for Sensor Deployment
@@
Ideally, the set of constructs that researchers want to investigate drives the choice of sensors. However, especially while estimating new measures, it can be challenging to ascertain a finite combination of sensors to deploy. Moreover, deployments are expensive and challenging. In longitudinal in-situ studies, even if issues like privacy concerns and maintenance of the sensing infrastructure can be mitigated, lack of participant compliance can deteriorate the quality of data [62]. Also, it may not be practical for all researchers to install a large suite of sensors and triage which signals are actually meaningful. In such cases, researchers can consider forms of low-burden sensing such as logging social media behavior [63, 64, 65, 66]. Therefore, we need to select sensors that sustain minimum burden on the user and still provide data consistently. Furthermore, the deployment of passive sensing technologies, as explored in a case study on just-in-time emotional support agents, underscores the importance of addressing boundary misalignments, data ownership, and power dynamics to facilitate successful implementation [67]. These considerations highlight the nuanced challenges of deploying personalized sensing systems in both work and nonwork contexts, emphasizing the need for careful alignment with user values and wellbeing definitions.
V-C Considerations for Machine Learning Models
The workplace is a dynamic context and therefore passive sensing studies tend to be in-situ studies over large periods of time. This approach comes with the caveat of missing data that can make it challenging to train models, especially when considering sensors that require constant skin contact that can be affected by movement or simply forgetfulness, as is the case with wearable devices. In addition, missing data is often not missing completely at random, meaning that simply ignoring missing data as is done in several studies can lead to a biased sample [68]. One way researchers can mitigate this for behavioral data is by exploiting complementary streams of data to impute the missing data [69, 20]. On the other hand, a large set of time-varying features opens an array of potential learning approaches, e.g., moment statistics, regularity, or auto–encoders. Researchers can empirically compare approaches to find the best model, but also consider other aspects like interpretability and transferability [70] to select learning approaches.
V-D Privacy and Ethical Considerations
@@
While we primarily highlighted the opportunities and potentials of these technologies and methodologies, these also come at a cost. For example, despite the best intentions, when instrumented at workplaces, these approaches can bear ethical and privacy concerns [71, 72, 73, 74, 75, 76]. There are lingering questions regarding the best practices of these algorithmic inferences for real-world measures and high-risk decisions because these approaches can be misused in ethically questionable ways. Consequently, these approaches and technologies can compromise privacy, defy expectations, and damage the trust between individuals and technologies [65]. Park et al.[77] identified key themes that negatively affect employee perception of AI use in the workplace, and suggest increased transparency and human-in-the-loop interventions as solutions. Table I shows that several approaches rely on sensing that occurs most often outside the workplace, e.g., sleep and physical activity, sensing for which there is no conceivable way to acquire consent from all involved parties such as speech activity, or sensing that could be carried out using the sensors in devices that could be provided by employers or fall into Bring-Your-Own-Device (BYOD) policies such as mobile phones. In fact, consenting to a passive sensing study in the workplace is complex because of the associated power dynamics [71, 78, 72]. A study by Chowdhary et al.[78] reveals that the passive sensing framework must support meaningful consent that is freely given, reversible, informed, enthusiastic, and specific.
@@
In fact, employer surveillance and employee’s subjective expectation of privacy share a competing relationship [79] and only a thin line of difference exists in perceiving the same technology as for surveillance or for supportive intervention [65]. There is a need to ensure that future work applications benefit individuals without sacrificing their privacy or forcing them to adopt behaviors solely on the need for increasing productivity in the workplace according to arbitrary metrics. A recent study by Das Swain et al.[72] interviewed 28 information workers about passive sensing in the workplace. Some concerns the workers talked about include the need for distinction between work and personal life, re-appraising sensor data through human evaluation, the dehumanizing nature of quantifying performance continuously, and selective sharing of data. This calls for the need to design guidelines that lead to responsible applications that balance the trade-offs of the risks and benefits associated with these data-driven human-centered assessments [80]. In addition, these guidelines should prevent the ingraining of biases in automated decisions and performance-evaluation systems, a danger made obvious by early attempts at developing AI tools for recruiting that resulted in biased results [81]. Considering the lack of diversity in the tech industry in the US [82], there is a considerable risk that the training of models based on existing employee data without special consideration of disproportionate representation could result in AI-based tools encoding biased behaviors that are not truly supporting task performance improvements for the companies or well-being benefits for the employees. These aspects should not be taken lightly and researchers should also incorporate metrics of AI fairness [83] and potential harms [71] when conducting future of work studies that predict performance outcomes.
V-E Limitations and Future Work
The existing body of work highlights several challenges with data-driven human-centered assessments to facilitate workplace outcomes. Many of these challenges suggest interesting future directions. In terms of operationalization, there is a lack of availability as well as a consensus about the definition of “ground-truth”. While self-reported measures are considered to be the gold-standard source of data, self-reported data bear their own limitations and biases. Moreover, from a methodological point of view, most of the approaches are not generalizable across problem scenarios or datasets — the same approach may not work on different populations or a different outcome measure. This necessitates accounting for domain information and problem-centric adjustments for model tuning.
Again, this field is evolving and most of the extant research is formative in revealing the feasibility and efficacy of these technologies. However, real-world and prospective adaptations of these approaches may bring in new challenges. For instance, the accuracy metrics tend to reflect the model performance at average, however, false positive on single instances (a single individual) may bear larger risks related to hiring or firing employees. This motivates thinking about novel measurement approaches that weigh in the risks and qualitative assessments beyond quantitative accuracy measures. These models not only need to be rigorously tested, but also proper guidelines need to be set about the role of algorithmic inference and human judgment. All in all, this literature survey shows us the potentials of passive and ubiquitous sensing methodologies within the context of Future of Work. However, there are several issues such as scalability, generalizability, privacy and ethical concerns which hinder the broad adoption of such technologies in spite of the widespread enthusiasm. Therefore, much work is needed to make workplace sensing ubiquitous. It is also important to highlight that our survey specifically focuses on recent advancements in the field, prioritizing works published since 2015. This criterion, among others, influenced the selection of studies, leading to the exclusion of older yet potentially relevant research. To ensure transparency and provide a broad perspective, we recommend readers refer to additional surveys that encompass the wider scope of developments in this area [84, 85, 86, 87, 88].
VI CONCLUSION

### v3

@@ 4.2 Measuring Workplace Performance
Driven by early successes in assessing organizational wellbeing, researchers have shifted to using passive sensing for workplace performance. By tracking behaviors, they aim to identify factors that boost productivity and develop interventions to improve it. Yet objectively and broadly quantifying workplace performance remains challenging. For example, while Amazon uses smartphone-based sensing to discourage drivers from checking their phones, frequent phone usage might be necessary in a consulting role. Therefore, researchers in the ubiquitous computing community have often turned to job performance inventories developed by psychologists as ground truth to measure perceived workplace performance across organizations and industries in a generalizable manner. Passive sensing is then used to predict inventory scores or categorize individuals into higher and lower performance tiers. Typically, these inventories measure task performance (e.g., Individual Task Proficiency [ITP], In-Role Behavior [IRB]) or behaviors fostering organizational effectiveness (e.g., Interpersonal and Organizational Deviance [IOD-ID, IOD-OD], Organizational Citizenship Behavior [OCB]).
@@
Task performance encompasses duties or actions formally acknowledged and rewarded by management [90]. Several studies have tried to predict ITP and IRB as measures of task performance. For instance, Feng et al.[32] study 84 full-time nurses over ten weeks, using Fitbit and Bluetooth proximity data to investigate how physiological responses vary across different hospital rooms. Combining a measure of “mutual dependency” between room use and physiological responses with aggregated features (e.g., min, mean, max, SD) yields adjusted $R^{2}$ scores of .11 for ITP and .12 for IRB. Another study[67] tracks 67 nurses wearing sociometric badges that measures physical activity, speech, face-to-face interaction, and proximity-based social networks. Multiple linear regression models using these features predict 49% of the variance in perceived workload and 63% of the variance in perceived productivity.
In studies of information workers, several works stand out. Das Swain et al.[21] track 249 individuals over 62 days using Bluetooth beacons to measure desk presence, work hours, and home time. They develop an organizational fit metric based on the alignment of individual routines with workplace norms, finding that routine fit, combined with Big Five Personality traits[74], explains up to 28% of the variance in IRB, with a significant positive association (p $<$ 0.05). In another 110-day study of 603 information workers, Das Swain et al.[24] introduce “organizational personas,” clustering individuals based on Big Five personality data and multimodal activity facets from a fitness tracker, phone agent (location, phone usage), and Bluetooth beacons (desk sessions). They find that activity facets contribute modest but statistically significant variance to ITP. Similarly, Pillai et al.[70] use an autoencoder-based approach to detect rare life events from multivariate mobile sensing data in 126 information workers. Mirjafari et al. [58] examine 554 information workers using a phone agent, beacons, and a Garmin fitness tracker. They apply K-Means clustering on ITP, IRB, OCB, IOD-ID, and IOD-OD scores to classify higher and lower workplace performers. Their findings show that features from fitness trackers alone achieve an AUROC of .72, phone agent features .65, and a combination of both .83, demonstrating the benefits of multimodal sensing in capturing workplace behaviors. In a follow-up study of 298 information workers, Mirjafari et al.[57] use autoencoder-generated features from mobile phones and a Garmin fitness tracker to predict daily job performance changes (improvement, decline, or no change) based on ITP, IRB, OCB, IOD-ID, and IOD-OD scores. Their XGBoost model achieves a 75% F1 score, significantly surpassing the 33% baseline. They also apply Gradient Analysis[84, 48] to provide user-actionable insights, helping individuals identify behaviors that influence performance. Similarly, Mark et al. [51] analyze email usage among 40 information workers, finding that increased time spent on emails correlates with lower self-reported productivity.
In the behavioral side of workplace performance inventories, IOD-ID and IOD-OD are measures of “bad” conduct in the workplace. Behaviors that indicate IOD-ID can involve cursing a co-worker, playing pranks, or making fun of someone. Behaviors that indicate IOD-OD can be tardiness or absenteeism, leaving work early without permission, putting little effort into work, among others [31]. On the other hand, OCB is a measure of “good” behavior. OCB is composed of behaviors that reflect altruism, conscientiousness, sportsmanship, courtesy and civic virtue [68]. Several studies have focused on these measures. Feng and Narayanan [34] show that behavioral consistency features from wearable data improve accuracy by 6% compared to summary features alone in a study of 97 hospital workers. Nepal et al. [64] analyze commute-based sensing data from 275 workers to predict OCB and IOD scores, achieving MAEs of less than 10%. In a previously mentioned study by Feng et al. [32], authors demonstrate that workplace physiological responses predict IOD-ID and IOD-OD (adjusted $R^{2}$ = 0.065 and 0.092). Multiple studies find associations between passive sensing data and workplace behaviors: Mirjafari et al. [58] with IOD and OCB, Das Swain et al. [24] with activity patterns, and beacon-based desk routines [21] with lower IOD, though accuracies remain moderate (ranging from 56.6% to 62.4%). The work by Nepal et al. [65] follows a different line of research. Instead of attempting to predict subjective performance measures from objective data, it intends to find what objective data is related to individual success in organizations, i.e., getting promoted. The authors note that if promotion can be detected and it produces objective physiological signals, researchers could learn from these signals and through the study of job promotion – which could be inferred – what behaviors could be behind the high performance that granted the promotion to measure them objectively, foster them and replicate them.
@@
Nurturing collaboration through open offices has become the norm in the modern era and workers face necessary interruptions throughout their workday to efficiently tackle problems. However, there are studies demonstrating that the nature and timing of interruptions negatively affect workplace productivity [6, 3] and increase worker stress [50]. Therefore, evaluating interruptibility has significant value to companies in terms of cost and time. To continuously measure a person’s interruptibility, Zuger et al. [97] study 13 software engineers over two weeks using wearable sensors in addition to keyboard and mouse interaction data. The ground truth in this study is collected using a pop-up question during work which asked the user to rate their interruptibility on a 1-7 Likert scale, and these ratings are further grouped into three states (12|345|67) for classification. Ultimately, they predicted interruptibility accurately (68.3%) by training a random forest on several features extracted from sensor modalities like heart rate, activity, and sleep.
Other works target efforts to support the design of future interventions in the workplace. Kimani et al. [46] develop a conversational agent that assists information workers with task management, scheduling, prioritization, and minimizing distractions. It leverages contextual data to determine optimal intervention times. In a six-day study with 24 participants, the agent analyzed speech, facial imagery, and computer usage, leading many users to become more aware of their work habits and adjust routines for improved productivity and wellbeing. Nepal et al. [63] introduce an LLM-powered productivity agent that utilizes computer-based telemetry to provide personalized assistance. They emphasize the importance of user-centric design, adaptability, and balancing personalization with privacy in AI-driven workplace tools. The integration of passive sensing with LLMs represents a major advancement, allowing for more contextually aware interventions, such as detecting when employees struggle with tasks or optimizing deep work periods. However, this convergence also raises concerns about data privacy, algorithmic bias, and the boundaries between AI assistance and workplace surveillance.
Kadoya et al.[42] use wearables to examine the link between emotional states and productivity, aiming to inform future interventions. Mark et al.[49] find that affect balance correlates with concentration and productivity, with sleep playing a key role. In another study, Schaule et al.[81] develop a smartwatch-based system to predict cognitive load and determine optimal moments for interruptions, while Di Lascio et al.[28] demonstrate how wearables can passively monitor workplace activity, supporting distraction management and break suggestions. Kucukozer-Cavdar et al.[47] create a model predicting office workers’ availability and break preferences using mobile sensing data. Das Swain et al.[20] assess computer-assisted protected (CAP) time in 89 information workers, showing significant benefits for those new to automated scheduling. Similarly, Saha et al. [77] further find that CAP enhances relaxation while reducing negative emotions like anger. These studies deepen our understanding of workplace performance and demonstrate ongoing efforts to develop productivity-enhancing interventions. With the rise of LLMs and AI, integrating passive sensing with advanced analytics presents a promising frontier, paving the way for more adaptive and intelligent workplace interventions.
5 DISCUSSION
@@
5.1 Considerations for Ground Truth
@@
5.2 Considerations for Sensor Deployment
@@
Ideally, the set of constructs that researchers want to investigate drives the choice of sensors. However, especially while estimating new measures, it can be challenging to ascertain a finite combination of sensors to deploy. Moreover, deployments are expensive and challenging. In longitudinal in-situ studies, even if issues like privacy concerns and maintenance of the sensing infrastructure can be mitigated, lack of participant compliance can deteriorate the quality of data [52]. Also, it may not be practical for all researchers to install a large suite of sensors and triage which signals are actually meaningful. In such cases, researchers can consider forms of low-burden sensing such as logging social media behavior [78, 25, 23, 27, 80]. Therefore, we need to select sensors that sustain minimum burden on the user and still provide data consistently. Furthermore, the deployment of passive sensing technologies, as explored in a case study on just-in-time emotional support agents, underscores the importance of addressing boundary misalignments, data ownership, and power dynamics to facilitate successful implementation [85]. These considerations highlight the nuanced challenges of deploying personalized sensing systems in both work and nonwork contexts, emphasizing the need for careful alignment with user values and wellbeing definitions.
5.3 Considerations for Machine Learning and AI at Work
The dynamic nature of workplaces makes passive sensing studies heavily reliant on longitudinal, in-situ data, often leading to missing data challenges. Sensors requiring constant skin contact, such as wearables, are particularly prone to gaps due to movement or non-compliance, and missing data is rarely random, making naïve omission approaches prone to bias [53]. Researchers can mitigate this by leveraging complementary data streams for imputation [79, 54]. Additionally, the high-dimensional, time-varying nature of workplace data demands careful model selection, with approaches such as moment statistics, regularity analysis, and auto-encoders offering varying trade-offs in performance, interpretability, and generalizability [93].
As AI and LLMs become more integrated into workplace sensing, they offer a powerful means of overcoming these challenges by extracting deeper insights from complex sensor data. LLMs, in particular, can enhance interpretability and actionability by translating raw sensing data into natural language explanations, making insights on wellbeing and productivity more accessible to both employees and managers. This ability to contextualize passive sensing data opens the door to more adaptive and personalized workplace interventions. However, the integration of AI in workplace monitoring also introduces ethical risks, including concerns about privacy, bias, and surveillance overreach. Ensuring fairness, transparency, and strong privacy protections is essential to building trust and preventing misuse. Future research should prioritize human-centered AI models that enhance workplace support while safeguarding worker autonomy and data security.
5.4 Privacy and Ethical Considerations
@@
While we primarily highlighted the opportunities and potentials of these technologies and methodologies, these also come at a cost. For example, despite the best intentions, when instrumented at workplaces, these approaches can bear ethical and privacy concerns [44, 19, 73, 15, 43, 4, 22]. There are lingering questions regarding the best practices of these algorithmic inferences for real-world measures and high-risk decisions because these approaches can be misused in ethically questionable ways. Consequently, these approaches and technologies can compromise privacy, defy expectations, and damage the trust between individuals and technologies [23]. Park et al.[69] identify key themes that negatively affect employee perception of AI use in the workplace, and suggest increased transparency and human-in-the-loop interventions as solutions. Table 1 shows that several approaches rely on sensing that occurs most often outside the workplace, e.g., sleep and physical activity, sensing for which there is no conceivable way to acquire consent from all involved parties such as speech activity, or sensing that could be carried out using the sensors in devices that could be provided by employers or fall into Bring-Your-Own-Device (BYOD) policies such as mobile phones. In fact, consenting to a passive sensing study in the workplace is complex because of the associated power dynamics [44, 12, 19]. A study by Chowdhary et al.[12] reveals that the passive sensing framework must support meaningful consent that is freely given, reversible, informed, enthusiastic, and specific.
In fact, employer surveillance and employee’s subjective expectation of privacy share a competing relationship [36], and only a thin line of difference exists in perceiving the same technology as for surveillance or for supportive intervention [23]. There is a need to ensure that future work applications benefit individuals without sacrificing their privacy or forcing them to adopt behaviors solely on the need for increasing productivity in the workplace according to arbitrary metrics. In fact, the awareness of being observed or “monitored” can lead individuals to deviate from their otherwise typical behaviors—or the phenomenon of observer effect or “Hawthorne” effect [76, 2]. Recent work by Das Swain et al.[18] investigates worker attitudes toward passive sensing systems, finding that acceptance varies significantly based on what is being sensed and how the data is used. Their work shows that workers are more accepting of systems that track digital time use or physical activity compared to visual or linguistic data, and prefer insights about mental wellbeing over performance metrics. Importantly, they find that workers resist automatic sharing of these insights with others, highlighting the need for worker agency in data sharing. In a different study, Das Swain et al.[19] conduct in-depth interviews with 28 information workers, revealing deeper concerns about workplace sensing. These include the need for clear boundaries between work and personal life, the importance of human interpretation of sensor data, apprehension about continuous performance quantification, and the desire for granular control over data sharing. Together, these studies emphasize the need for guidelines that lead to responsible applications that balance the trade-offs of risks and benefits associated with these data-driven human-centered assessments [39].
@@
In addition, these guidelines should prevent the ingraining of biases in automated decisions and performance-evaluation systems, a danger made obvious by early attempts at developing AI tools for recruiting that resulted in biased results [26]. Considering the lack of diversity in the tech industry in the US [1], there is a considerable risk that the training of models based on existing employee data without special consideration of disproportionate representation could result in AI-based tools encoding biased behaviors that are not truly supporting task performance improvements for the companies or well-being benefits for the employees. These aspects should not be taken lightly and researchers should also incorporate metrics of AI fairness [96] and potential harms [44] when conducting future of work studies that predict performance outcomes.
5.5 Limitations and Future Work
The existing literature reveals multiple challenges in using data-driven, human-centered assessments for workplace outcomes. Conflicting definitions of “ground-truth” and limited availability of reliable data pose significant issues, while self-reported measures can introduce biases. Methodologically, many approaches lack generalizability across different populations or outcome measures, requiring domain-specific adjustments and tuning. This field is still evolving, and current studies are largely formative in demonstrating feasibility. Yet real-world applications may introduce new complications. For instance, average accuracy metrics can mask high-stakes risks for individuals in hiring or firing decisions, underscoring the need for novel measurement strategies that consider both risk and qualitative factors. Ensuring rigorous testing of these models—and establishing clear guidelines around algorithmic inference and human judgment—is essential.
Overall, this survey highlights the potential of passive, ubiquitous sensing in the Future of Work, but challenges such as scalability, privacy, and ethical concerns still limit widespread adoption. More research is needed to overcome these barriers and make workplace sensing truly ubiquitous. Our review specifically emphasizes studies published since 2015, which led to the exclusion of older but potentially relevant work. For a broader perspective, we recommend additional surveys  [95, 61, 89, 71, 7]. Finally, rapid advances in AI and LLMs offer promising avenues for enhancing workplace interventions via passive sensing. Future work should integrate these technologies to yield interpretable, privacy-preserving insights while upholding fairness and transparency.
6 Conclusion